OpenStudy (anonymous):

I need help to check this equation for extraneous solutions. 4|3x+4|=4x+8 I need the steps to if possible

OpenStudy (anonymous):

\[\left| 3x+4 \right|=3x+4 if 3x+4\ge 0, x \ge \frac{ -4 }{ 3 }\] \[\left| 3x+4 \right| <0 if x < \frac{ -4 }{3 },annd \left| 3x+4 \right| =-\left( 3x+4 \right)\] Now you can solve botth the equations.